OpenStudy (anonymous):

$1800 is deposited in an account that pays 4.5% annual interest, compounded continuously. when will the balance be $5000

OpenStudy (anonymous):

When you get a job and stop depending on interest

OpenStudy (anonymous):

@UnkleRhaukus haven't heard from you all day :( advoiding me much??

OpenStudy (anonymous):

hahah @DylanH

OpenStudy (anonymous):

lol but for the answer do you want that in years or what?

OpenStudy (anonymous):

well i want you to help me solve dont give me the answer please.. and i dont think it matters

OpenStudy (anonymous):

Its a repeating decimal

OpenStudy (unklerhaukus):

i have been working on my own problems

OpenStudy (anonymous):

So first you find out what you get each year

OpenStudy (anonymous):

which is 1800 x .045

OpenStudy (anonymous):

@DylanH got it

OpenStudy (anonymous):

is there not a formula???

OpenStudy (anonymous):

no you had to do it step by step

OpenStudy (anonymous):

@UnkleRhaukus Oh i understand!

OpenStudy (anonymous):

@satellite73 WHATCHU GET?!?!?

OpenStudy (anonymous):

first off lets to this problem, then we do yours

OpenStudy (anonymous):

81?

OpenStudy (anonymous):

then after you know the annual rate of 81$ you subtract 1800 from 5000 to see how much you need

OpenStudy (anonymous):

3200

OpenStudy (anonymous):

Nice! Now you divide that number by 81 to see how long it will take

OpenStudy (anonymous):

\[5000=1800e^{.045t}\] divide by1800 take the log divide by .045

OpenStudy (anonymous):

39.51

OpenStudy (anonymous):

thats how many yard it will take

OpenStudy (anonymous):

*years autocorrect

OpenStudy (anonymous):

\[\frac{25}{9}=e^{.045t}\] \[\ln(\frac{25}{9})=.045t\] \[t=\ln(\frac{25}{9})\div.045\]

OpenStudy (anonymous):

Or you could do it the hard way^^

OpenStudy (anonymous):

about 22.7 years no other way to do it now on to @DylanH question i will write it out there when you are ready

OpenStudy (anonymous):

im ready

OpenStudy (anonymous):

wait why do i have 2 different answers....:/

OpenStudy (anonymous):

@satellite73 @DylanH

OpenStudy (anonymous):

sate gave me 22.7 and Dyl we got 39.51